OpenStudy (anonymous):

I need help and a lot of it solve radical equation x: square root X+2= 2

OpenStudy (lgbasallote):

\[\sqrt x + 2 = 2??\]

OpenStudy (anonymous):

Square both sides.

OpenStudy (anonymous):

reprat the question properly

OpenStudy (anonymous):

presently it gives you, x=0

OpenStudy (anonymous):

Yes lgbasallote that is how it is

OpenStudy (lgbasallote):

just subtract 2 to both sides..what will you have?

OpenStudy (anonymous):

i do not know how to do the square root symbol on the computer

OpenStudy (lgbasallote):

put \sqrt in your equations editor

OpenStudy (anonymous):

O cant be the answer cause it is multple choice and its not listed 6,4,2,16

OpenStudy (lgbasallote):

well could it be \[\sqrt{x + 2} = 2\] that will make more sense

OpenStudy (anonymous):

yes that is the equation \[X: \sqrt{x +2 =2}\]

OpenStudy (anonymous):

then, it will be 16 \[(\sqrt{x}+2)^{2} = 2^{2}\] or, \[x+ 4\sqrt{x} +4 = 4\] or, \[x+ 4\sqrt{x} = 0\] or, \[x = -4\sqrt{x}\] or, \[x ^{2} = 16x\] or, \[x = 16\]

OpenStudy (anonymous):

in second case: \[\sqrt{x+2} = 2\] or, \[x+2 = 4\] (by square both side) or, x=2

OpenStudy (anonymous):

okay thanks I have no idea how to do this stuff will you all please help me

OpenStudy (anonymous):

\[X: \sqrt{2x = 5}\]

OpenStudy (anonymous):

Im totally lost